A basketball player drops a 0.60-kg basketball vertically so that it is traveling at 6.3 m/s when it reaches the floor. The ball rebounds upward at a speed of 4.7 m/s.

B.) Determine the average net force that the floor exerts on the ball if the collision lasts 0.12 s.

Express your answer to two significant figures and include the appropriate units.

NSonB =
Answer #1

Average force

F = m ( vf - vi ) t

= ( 0.6 kg ) (6.3 m/s - 4.7 m/s ) / 0.12

=8 kg .m/s^2
Answer #2

applying newton's second law,

F=m(v-u)/dt=0.6(6.3-(-4.7))/0.12

=50 newton